# Neural Approaches for Multi-Objective Routing on Multigraphs

### v2

@@ B.3 MatNet-Based Benchmark Model
Lastly, we note that MatNet is not well-defined on multigraphs due to the fact that it utilizes the distance matrix to represent unique distances between nodes. It is possible to cast multigraphs into simple graphs, but not without increasing the number of nodes and edges drastically. Thus, we must pre-prune a multigraph before passing it through the encoder. This can be done with some sort of heuristic in the general case, but under linear scalarization and considering the MOTSP, an optimal pruning method is to remove edges with suboptimal scalarized cost. We utilize this pruning method for both multigraph problems. Moreover, it can be noted that the encoder needs to be rerun for each new preference, which results in a slower architecture than if the encoding is done once.
@@
Figure 3: MatNet-Based Model (MBM)
Appendix C Model Training
@@
C.1 Curriculum Learning Details
@@ C.2 Training Times
00505050501001001001001501501501502002002002009090909095959595100100100100EpochHV ratio (%)MOTSP TMAT2000505050501001001001001501501501502002002002009696969698989898100100100100EpochMGMOTSP FLEX2-20MBMGMS-EBGMS-DH
@@
00505050501001001001001501501501502002002002009090909095959595100100100100EpochHV ratio (%)MOTSP TMAT500050505050100100100100150150150150200200200200949494949696969698989898100100100100EpochMGMOTSP FLEX2-50GMS-EBGMS-DHInstance Size Switch
Figure 5: HV as percentage of best performance for GMS during training. It is enough to train on 20 nodes to obtain a majority ($\approx 98\%$) of the performance. After 5 epochs of training on instances with 50 nodes, the models reach $>99.5\%$ of their final performance.
Appendix D Routing Problems and Distributions
@@ D.1 Distributions
  • TMAT - Asymmetric distance matrix obeying the triangle inequality (Kwon et al. 2021). The distance matrix is first populated with random independent elements and then post-processed until the triangle inequality is satisfied. For a problem with $m$ distances, we sample $m$ matrices independently.
  • XASY - Extremely asymmetric distance matrix (Gaile et al. 2022). The distance matrix does not satisfy the triangle inequality. It is obtained by independently sampling each element uniformly between 0 and 1. Again, we sample all $m$ matrices independently.
@@
[…]
D.2 MOTSP
Problem Description
@@ D.2 MOTSP
In the multi-objective traveling salesman problem the goal is to visit all nodes/customers and return to the starting position. The Euclidean version of this problem has been widely treated in the learning-for-routing community, e.g., by Lin, Yang, and Zhang (2022) and Fan et al. (2025). In our case, we sample two distances, corresponding to the two objectives, between each pair of customers. The goal is to design a tour that minimizes the sum of the distances. A tour can start from any node.
Data Generation
@@
Edge distances are sampled according to the EUC, TMAT and XASY distributions above.
Model Details
@@
There are no node features for this problem, hence both the GREAT-based and MatNet-based encoders have the two distances as edge features, while the latter uses zero-vectors as ”row” input.
Evaluation Details
@@
For all distributions, we utilize $(15,15)$, $(30,30)$ and $(60,60)$ as reference for HV calculation for $20$, $50$ and $100$ nodes respectively. In the zero-shot experiments, we have $(90,90)$ as reference for 150 nodes and $(120,120)$ for 200 nodes.
D.3 MOCVRP
Problem Description
@@ D.3 MOCVRP
In the multi-objective capacitated vehicle routing problem, a truck starts at a depot and must pick up goods from a specified number of customers. The vehicle has a fixed capacity while each customer has a demand. When the capacity is full, the vehicle must return to the depot. As many vehicles as required can be utilized to meet the demand of all customers.
@@
Similarly to the TSP, this problem has been widely treated in the Euclidean setting using learning-based methods. As is common in the multi-objective setting (Lin, Yang, and Zhang 2022), we assume one objective is the traveled distance, while the other is the longest tour for a single vehicle (makespan). As such, each node pair is only associated with one distance.
Data Generation
@@
We assume that the capacity of the truck is 30, 40 and 50 for 20, 50 and 100 nodes respectively, while the demands are sampled uniformly from the set $\{1,\dots,9\}$. This is the same setting as in e.g., Kool, van Hoof, and Welling (2018) and Lin, Yang, and Zhang (2022). Edge distances are sampled according to the EUC, TMAT and XASY distributions.
Model Details
@@
In the decoding, we augment the routing context (represented by the query $q_{t}$) with the current load of the vehicle. For the GREAT-based encoders, we let the features of an edge be its distance as well as the demand of the customer at the end. On the other hand, the MatNet-based encoders only utilize the distance as an edge-feature while the demand is used as a node feature.
Evaluation Details
@@
For all distributions, we utilize $(15,3)$, $(40,3)$ and $(60,3)$ as reference for HV calculation for the three considered instance sizes.
D.4 MGMOTSP
Problem Description
@@ D.4 MGMOTSP
As for the MOTSP, the goal of the multigraph MOTSP is to visit all nodes and return to the start. The only difference is that the underlying structure is a multigraph. Similarly, we define two distances for each node pair and these are associated with the objectives of the problem.
@@
While this variant of the TSP on a multigraph has not been treated before to our knowledge (neither using learning nor other methods), some stochastic single-objective variants have been proposed by e.g., Tadei, Perboli, and Perfetti (2017) and Maggioni, Perboli, and Tadei (2014). Moreover, we believe that the MGMOTSP is a suitable benchmark problem for our solvers, as it is simple and easy to find good baselines for.
Data Generation
[…]
Model Details
@@
As the MOTSP, the MGMOTSP lacks node features. Hence, we utilize the same setup for the models as in the MOTSP. Moreover, for the MatNet-based model, we sparsify the underlying multigraph given each preference before it is fed into the encoder. This is done by removing all edges with suboptimal linearly scalarized cost compared to a parallel edge. To make this pruning strategy exact, we train MBM using linear reward instead of the Chebyshev reward.
Evaluation Details
[…]
D.5 MGMOTSPTW
Problem Description
@@ D.5 MGMOTSPTW
The goal in MGMOTSPTW is to create a tour that visits all customers and returns to the starting position. However, in this case we assume there is a fixed depot and that each customer/node is associated with a time-window. Preferably, the vehicle should arrive within the time-window for each customer. As such, we let one objective be the number of violated time-windows while the other is the traveled distance. Accordingly, each edge connecting two locations is associated with a travel time and a distance.
@@
Multi-objective problems with similar objectives are fairly widespread in literature (Legrand et al. 2025), although not on multigraphs. In the multigraph framework, this problem is reminiscent of the vehicle routing problem with time-windows of Ben Ticha et al. (2017), but with a second objective that is the soft variant of the time-window constraint.
Data Generation
[…]
@@
Model Details
In the decoder, we enhance the routing context with an embedding of the current time. In the encoder, the node features in this case are the starting time and end time of the service, while the edge features are distance and travel time. Consequently, we augment the input features of GREAT so that an edge is associated with its distance, travel time and the time-window of the ending node. As usual for MatNet, the node features are used as ”row” inputs while the edge features are used in the distance matrix.
@@
Evaluation Details
In this case, we let the reference value corresponding to the time-window violation be $N+5$, while the reference value corresponding to the distance is the same as in the MGMOTSP case.
Appendix E Setup for MOEAs
As a base for the implementation of NSGA-II, NSGA-III and MOEA/D, we utilized Pymoo [^2] https://pymoo.org/ (Blank and Deb 2020) and as a base for MOGLS, we utilized GeneticLocalSearch_TSP [^3] https://github.com/kevin031060/Genetic˙Local˙Search˙TSP (Li, Zhang, and Wang 2021). For the former three, we limit the number of iterations to $1000$ while the latter is limited to $10000$. Additionally, local search strategies are a necessity for efficient usage of genetic algorithms in combinatorial optimization. For the MOTSP, we utilized the 2-opt heuristic of Jaszkiewicz (2002), while we utilized the problem-specific local search and chromosome decoding strategy of Lacomme, Prins, and Sevaux (2006) for the MOCVRP.
However, in both cases we find that by accepting the best local move instead of just an improving one, the performance on large asymmetric problems improves substantially. As this affects the runtime negatively (due to the need to search through all moves), we limit the number of local moves in each search depending on the problem size. Additionally, for the algorithms not requiring preference weights for each individual (NSGA-II and NSGA-III), we must calculate such weights to compare the improvements. In the bi-objective case, this is done using the strategy of Lacomme, Prins, and Sevaux (2006) according to
@@ Appendix E Setup for MOEAs
In the multigraph setting, we design our own setup to tailor the MOEAs. To start with, inspired by one representation from Beke, Weiszer, and Chen (2021) regarding shortest path problems, we let a chromosome for a tour be given by a permutation of the nodes multiplied with $100$ added with the index of the edge to the next node. For example, in an MGMOTSP with 2 parallel edges and 5 nodes, a chromosome might be
@@
Appendix F Ablation Study
@@
To ensure a fair comparison, we evaluate both learned pruning and simple pruning using linearly scalarized rewards and Chebyshev rewards. In the linear case, simple pruning is optimal by Proposition 1. Moreover, the learned pruning model is trained using $K_{1}=4$ edge selections per instance. Accordingly, we use 4$\times$ more POMO samples when training the model with simple pruning to match the training budget.
@@
Appendix G Complementary Numerical Results
In this section, we show additional numerical results for more distributions, benchmarks and instance sizes.
G.1 Problems on Simple Graphs
Table 6 and 7 show results for the MOTSP and MOCVRP respectively. Apart from the benchmarks in the main text, we present the performance of three simple greedy heuristics on the MOTSP: Nearest Neighbor (NN), Nearest Insertion (NI) and Farthest Insertion (FI). Like LKH and OR-tools, these utilize the linear scalarization to transform the MO problem into single-objective subproblems. Additionally, we show results on Euclidean instances (EUC). For these, there are many existing models in literature and thus we replace MBM with the Conditional Neural Heuristic (CNH) of Fan et al. (2025). This model is augmented with the Euclidean augmentation of Lin, Yang, and Zhang (2022).
On the EUC distribution, our models are generally slightly worse than CNH. This is to be expected though, as a common pattern in previous studies is that architectures built for directed graph inputs perform worse on EUC than those specialized for coordinate inputs (Lischka et al. 2025; Kwon et al. 2021).
@@ G.1 Problems on Simple Graphs
G.2 Multigraph Problems
We display additional results for the MGMOTSP and MGMOTSPTW in Tables 8-11. As for the MOTSP, we include NN, NI and FI for MGMOTSP. In the case of MGMOTSPTW, we also include results for LKH and OR-tools when treating the problem as the MGMOTSP. That is, we neglect the time-windows and treat the time duration of each edge exactly as the distance. We call these methods implicit LKH and implicit OR-tools, as they implicitly minimize the time-window violations by reducing the time the tour takes. Additionally, for the MGMOTSP, we show results for FLEX5555 and FIX5555, i.e., when there are 5 parallel edges between each node pair.
@@ G.2 Multigraph Problems
[…]
@@
[…]
FLEX2-20 FLEX2-50
HV Gap Time HV Gap Time
Implicit LKH 0.27 66.74% (1.1m) 0.14 84.11% (6.7m)
Implicit OR Tools 0.24 70.78% (3.4m) 0.13 85.34% (25m)
MOGLS 0.64 22.92% (2.7h) 0.35 60.96% (21h)
NSGA-II 0.69 16.55% (2.4h) 0.51 43.47% (35h)
NSGA-III 0.69 16.01% (2.5h) 0.50 44.53% (36h)
MOEA/D 0.75 9.19% (2.7h) 0.60 32.79% (34h)
MBM 0.72 12.89% (4.9s) 0.80 11.09% (14s)
MBM (aug) 0.73 11.98% (34s) 0.81 10.10% (1.8m)
GMS-EB 0.81 1.95% (4.3s) 0.89 0.97% (60s)
GMS-DH 0.80 3.37% (3.3s) 0.85 5.10% (13s)
GMS-EB (aug) 0.82 0.00% (31s) 0.90 0.00% (7.9m)
GMS-DH (aug) 0.82 0.33% (20s) 0.87 2.69% (1.6m)
GMS-DH Simple 0.76 8.18% (3.3s) 0.83 7.01% (13s)
[…]
FIX2-20 FIX2-50
HV Gap Time HV Gap Time
Implicit LKH 0.28 67.24% (1.3m) 0.15 84.24% (6.5m)
Implicit OR Tools 0.25 71.02% (3.4m) 0.14 85.49% (26m)
MOGLS 0.62 28.85% (2.8h) 0.41 55.72% (21h)
NSGA-II 0.73 15.93% (2.7h) 0.56 39.66% (35h)
NSGA-III 0.71 18.88% (2.5h) 0.53 43.42% (35h)
MOEA/D 0.74 14.99% (2.8h) 0.60 35.73% (36h)
MBM 0.53 38.64% (4.9s) 0.64 31.18% (14s)
MBM (aug) 0.54 38.10% (35s) 0.66 28.58% (1.9m)
GMS-EB 0.86 1.17% (4.3s) 0.93 0.55% (59s)
GMS-DH 0.85 2.47% (3.4s) 0.91 1.74% (12s)
GMS-EB (aug) 0.87 0.00% (31s) 0.93 0.00% (7.9m)
GMS-DH (aug) 0.87 0.46% (20s) 0.92 0.78% (1.7m)
GMS-DH Simple 0.76 12.07% (3.2s) 0.85 8.93% (12s)
[…]
G.3 Generalization to Larger Instances
In Tables 12-14, we report more generalization results for our methods in the MOTSP and MGMOTSP cases. Notably, even when introducing significantly more edges by testing on FLEX10101010 and FIX10101010, our methods remain competitive zero-shot.

### v4

@@ C.3 MatNet-Based Benchmark Model
Figure 4: MatNet-Based Model (MBM).
Lastly, we note that MatNet is not well-defined on multigraphs due to the fact that the distance matrices $D_{1},\dots,D_{f}$ represent unique edge features between nodes. Thus, we must pre-prune a multigraph before passing it through the encoder. This can be done with some sort of heuristic in the general case, but under linear scalarization and considering the MOTSP, an optimal pruning method is to remove edges with suboptimal scalarized cost. We utilize this pruning method for both multigraph problems. It can then be noted that the encoder needs to be rerun for each new preference, which results in a slower architecture than if the encoding is done once.
Appendix D Model Training
@@
D.1 Curriculum Learning Details
@@ D.2 Training Times
0505010010015015020020090909595100100EpochHV ratio (%)MOTSP TMAT200505010010015015020020096969898100100EpochMGMOTSP FLEX2-20MBMGMS-EBGMS-DH
@@
0505010010015015020020090909595100100EpochHV ratio (%)MOTSP TMAT5005050100100150150200200949496969898100100EpochMGMOTSP FLEX2-50GMS-EBGMS-DHInstance Size Switch
Figure 6: HV as percentage of best performance for GMS during training. It is enough to train on 20 nodes to obtain a majority ($\approx 98\%$) of the performance. After 5 epochs of training on instances with 50 nodes, the models reach $99.5\%$ of their final performance.
Appendix E Routing Problems and Distributions
@@ E.1 Distributions
  • TMAT - Asymmetric distance matrix obeying the triangle inequality (Kwon et al., 2021). The distance matrix is first populated with random independent elements and then post-processed until the triangle inequality is satisfied. For a problem with $m$ distances, we sample $m$ matrices independently.
  • XASY - Extremely asymmetric distance matrix (Gaile et al., 2022). The distance matrix does not satisfy the triangle inequality. It is obtained by independently sampling each element uniformly between 0 and 1. Again, we sample all $m$ matrices independently.
@@
  • FLEXxx - Flexible number of edges between each node pair. Here we sample $x$ vectors of dimension $m$ representing edge attributes for edges between a node pair. Each attribute for each edge is sampled uniformly between 0 and 1. Then we remove dominated edges until only a Pareto set of edges remains.
  • FIXxx - Fixed number of edges between each node pair. We start by sampling $x$ vectors in the same way as the FLEX-distribution. Then we sort each attribute so that no edge is dominated. For $m=2$, we sort one attribute in descending order and one attribute in ascending order. For $m=3$ (showed in Appendix H), we sort one attribute in descending order, one in ascending order and keep one attribute in random order.
Note that neither FLEXxx nor FIXxx obeys the triangle inequality, although such an assumption holds for real instances. We highlight that these distributions are utilized since they are simple and easy to sample from, creating a more transparent picture of the model performance. Motivated by the fact that XASY seems more difficult to learn than TMAT, we hypothesize that the triangle inequality would contribute positively to model performance.
E.2 MOTSP
@@
Problem Description. In the multi-objective traveling salesman problem the goal is to visit all nodes/customers and return to the starting position. The Euclidean version of this problem has been widely treated in the learning-for-routing community, e.g., by Lin et al. (2022) and Fan et al. (2025). In our case, we sample $m$ distances, corresponding to the $m$ objectives, between each pair of customers. The goal is to design a tour that minimizes the sum of the distances. A tour can start from any node.
@@
Data Generation. Edge distances are sampled according to the EUC, TMAT and XASY distributions above.
@@
Model Details. There are no node features for this problem, hence both the GREAT-based and MatNet-based encoders have the two distances as edge features, while the latter uses zero-vectors as ”row” input.
@@
Evaluation Details. For all distributions in the bi-objective case, we utilize $(15,15)$, $(30,30)$ and $(60,60)$ as reference for HV calculation for $20$, $50$ and $100$ nodes respectively. In the zero-shot experiments, we have $(90,90)$ as reference for 150 nodes and $(120,120)$ for 200 nodes. In the tri-objective case, we use $(15,15,15)$. $(30,30,30)$ and $(60,60,60)$ for $20$, $50$ and $100$ nodes respectively.
E.3 MOCVRP
@@
Problem Description. In the multi-objective capacitated vehicle routing problem, a truck starts at a depot and must pick up goods from a specified number of customers. The vehicle has a fixed capacity while each customer has a demand. When the capacity is full, the vehicle must return to the depot. As many vehicles as required can be utilized to meet the demand of all customers.
@@
Similarly to the TSP, this problem has been widely treated in the Euclidean setting using learning-based methods. As is common in the multi-objective setting (Lin et al., 2022), we assume one objective is the traveled distance, while the other is the longest tour for a single vehicle (makespan). As such, each node pair is only associated with one distance.
@@
Data Generation. We assume that the capacity of the truck is 30, 40 and 50 for 20, 50 and 100 nodes respectively, while the demands are sampled uniformly from the set $\{1,\dots,9\}$. This is the same setting as in e.g., Kool et al. (2018) and Lin et al. (2022). Edge distances are sampled according to the EUC, TMAT and XASY distributions.
@@
Model Details. In the decoding, we augment the routing context (represented by the query $q_{t}$) with the current load of the vehicle. For the GREAT-based encoders, we let the features of an edge be its distance as well as the demand of the customer at the end. On the other hand, the MatNet-based encoders only utilize the distance as an edge-feature while the demand is used as a node feature.
@@
Evaluation Details. For all distributions, we utilize $(15,3)$, $(40,3)$ and $(60,3)$ as reference for HV calculation for the three considered instance sizes.
E.4 MGMOTSP
@@
Problem Description. As for the MOTSP, the goal of the multigraph MOTSP is to visit all nodes and return to the start. The only difference is that the underlying structure is a multigraph. Similarly, we define $m$ distances for each node pair and these are associated with the objectives of the problem.
@@
While this variant of the TSP on a multigraph has not been treated before to our knowledge (neither using learning nor other methods), some stochastic single-objective variants have been proposed by e.g., Tadei et al. (2017) and Maggioni et al. (2014). Moreover, we believe that the MGMOTSP is a suitable benchmark problem for our solvers, as it is simple and easy to find good baselines for.
Data Generation. We sample distances between the nodes according to the FLEXxx and FIXxx distributions.
@@
Model Details. As the MOTSP, the MGMOTSP lacks node features. Hence, we utilize the same setup for the models as in the MOTSP. Moreover, for the MatNet-based model, we sparsify the underlying multigraph given each preference before it is fed into the encoder. This is done by removing all edges with suboptimal linearly scalarized cost compared to a parallel edge. To make this pruning strategy exact, we train MBM using linear reward instead of the Chebyshev reward.
Evaluation Details. For the FLEXxx distribution in the bi-objective case, we utilize $(15,15)$, $(30,30)$ and $(60,60)$ as reference for instance sizes 20, 50 and 100, regardless of $x$. In the zero-shot experiments we set the references to $(90,90)$ and $(120,120)$ for FLEXxx with 150 and 200 nodes. For FIXxx, the reference is given by $(N,N)$, where $N$ is the node count. When considering three objectives, we use $(15,15,15)$, $(30,30,30)$ and $(60,60,60)$ for FLEXxx and $(N,N,N)$ for FIXxx.
E.5 MGMOTSPTW
@@
Problem Description. The goal in MGMOTSPTW is to create a tour that visits all customers and returns to the starting position. However, in this case we assume there is a fixed depot and that each customer/node is associated with a time-window. Preferably, the vehicle should arrive within the time-window for each customer. As such, we let one objective be the number of violated time-windows while the other is the traveled distance. Accordingly, each edge connecting two locations is associated with a travel time and a distance.
@@
Multi-objective problems with similar objectives are fairly widespread in literature (Legrand et al., 2025), although not on multigraphs. In the multigraph framework, this problem is reminiscent of the vehicle routing problem with time-windows of Ben Ticha et al. (2017), but with a second objective that is the soft variant of the time-window constraint.
Data Generation. Distances and travel times are sampled according to the FLEXxx and FIXxx distributions. Additionally, we sample time-windows according to the ”medium” distribution of Chen et al. (2024) and Bi et al. (2024). In this setting, time-windows are sampled uniformly and independently, which results in difficult instances where there is a true trade-off between optimal paths and paths satisfying the time-window constraint. We also argue that this premise is simple for initial evaluation of our models, especially in comparison to e.g., the ”hard” instances of Chen et al. (2024), which test out-of-distribution generalization.
@@
Model Details. In the decoder, we enhance the routing context with an embedding of the current time. In the encoder, the node features in this case are the starting time and end time of the service, while the edge features are distance and travel time. Consequently, we augment the input features of GREAT so that an edge is associated with its distance, travel time and the time-window of the ending node. As usual for MatNet, the node features are used as ”row” inputs while the edge features are used in the distance matrix.
@@
Evaluation Details. In this case, we let the reference value corresponding to the time-window violation be $N+5$, while the reference value corresponding to the distance is the same as in the MGMOTSP case.
Appendix F Setup for MOEAs
As a base for the implementation of NSGA-II, NSGA-III and MOEA/D, we utilized Pymoo [^2] https://pymoo.org/ (Blank & Deb, 2020) and as a base for MOGLS, we utilized GeneticLocalSearch_TSP [^3] https://github.com/kevin031060/Genetic_Local_Search_TSP (Li et al., 2021). For the former three, we limit the number of iterations to $1000$ while the latter is limited to $10000$. Additionally, local search strategies are a necessity for efficient usage of genetic algorithms in combinatorial optimization. For the MOTSP, we utilized the 2-opt heuristic of Jaszkiewicz (2002), while we utilized the problem-specific local search and chromosome decoding strategy of Lacomme et al. (2006) for the MOCVRP.
However, in both cases we find that by accepting the best local move instead of just an improving one, the performance on large asymmetric problems improves substantially. As this affects the runtime negatively (due to the need to search through all moves), we limit the number of local moves in each search depending on the problem size. Additionally, for the algorithms not requiring preference weights for each individual (NSGA-II and NSGA-III), we must calculate such weights to compare the improvements. In the bi-objective case, this is done using the strategy of Lacomme et al. (2006) according to
@@ Appendix F Setup for MOEAs
In the multigraph setting, we design our own setup to tailor the MOEAs. To start with, inspired by one representation from Beke et al. (2021) regarding shortest path problems, we let a chromosome for a tour be given by a permutation of the nodes multiplied with $100$ added with the index of the edge to the next node. For example, in an MGMOTSP with 2 parallel edges and 5 nodes, a chromosome might be
@@
Appendix G Ablation Study
@@
[…]
@@
Appendix H Complementary Numerical Results
In this section, we show additional numerical results for more problems, distributions, benchmarks and instance sizes.
H.1 Problems on Simple Graphs
Table 6 and 7 show results for the MOTSP and MOCVRP respectively. Apart from the benchmarks in the main text, we present the performance of three simple greedy heuristics on the MOTSP: Nearest Neighbor (NN), Nearest Insertion (NI) and Farthest Insertion (FI). Like LKH and OR-tools, these utilize the linear scalarization to transform the MO problem into single-objective subproblems. Additionally, we show results on Euclidean instances (EUC). For these, there are many existing models in literature and thus we replace MBM with the Conditional Neural Heuristic (CNH) of Fan et al. (2025). This model is augmented with the Euclidean augmentation of Lin et al. (2022).
On the EUC distribution, our models are generally slightly worse than CNH. This is to be expected though, as a common pattern in previous studies is that architectures built for directed graph inputs perform worse on EUC than those specialized for coordinate inputs (Lischka et al., 2025; Kwon et al., 2021).
@@ H.1 Problems on Simple Graphs
H.2 Multigraph Problems
We display additional results for the MGMOTSP and MGMOTSPTW in Tables 8-10. As for the MOTSP, we include NN, NI and FI for MGMOTSP. In the case of MGMOTSPTW, we also include results for LKH and OR-tools when treating the problem as the MGMOTSP. That is, we neglect the time-windows and treat the time duration of each edge exactly as the distance. We call these methods implicit LKH and implicit OR-tools, as they implicitly minimize the time-window violations by reducing the time the tour takes. Additionally, for the MGMOTSP, we show results for FLEX55 and FIX55, i.e., when there are 5 parallel edges between each node pair.
Table 8: MGMOTSP on FLEXxx with different instances sizes (20, 50 and 100) and number of edges (2, 5).
@@ H.2 Multigraph Problems
Table 9: MGMOTSP on FIXxx with different instances sizes (20, 50 and 100) and number of edges (2, 5).
@@
Table 10: MGMOTSPTW on FLEXxx and FIXxx with different instances sizes (20, 50).
FLEX2-20 FLEX2-50 FIX2-20 FIX2-50
HV Gap Time HV Gap Time HV Gap Time HV Gap Time
Implicit LKH 0.27 66.74% (1.1m) 0.14 84.11% (6.7m) 0.28 67.24% (1.3m) 0.15 84.24% (6.5m)
Implicit OR Tools 0.24 70.78% (3.4m) 0.13 85.34% (25m) 0.25 71.02% (3.4m) 0.14 85.49% (26m)
MOGLS 0.64 22.92% (2.7h) 0.35 60.96% (21h) 0.62 28.85% (2.8h) 0.41 55.72% (21h)
NSGA-II 0.69 16.55% (2.4h) 0.51 43.47% (35h) 0.73 15.93% (2.7h) 0.56 39.66% (35h)
NSGA-III 0.69 16.01% (2.5h) 0.50 44.53% (36h) 0.71 18.88% (2.5h) 0.53 43.42% (35h)
MOEA/D 0.75 9.19% (2.7h) 0.60 32.79% (34h) 0.74 14.99% (2.8h) 0.60 35.73% (36h)
MBM 0.72 12.89% (4.9s) 0.80 11.09% (14s) 0.53 38.64% (4.9s) 0.64 31.18% (14s)
MBM (aug) 0.73 11.98% (34s) 0.81 10.10% (1.8m) 0.54 38.10% (35s) 0.66 28.58% (1.9m)
GMS-EB 0.81 1.95% (4.3s) 0.89 0.97% (60s) 0.86 1.17% (4.3s) 0.93 0.55% (59s)
GMS-DH 0.80 3.37% (3.3s) 0.85 5.10% (13s) 0.85 2.47% (3.4s) 0.91 1.74% (12s)
GMS-EB (aug) 0.82 0.00% (31s) 0.90 0.00% (7.9m) 0.87 0.00% (31s) 0.93 0.00% (7.9m)
GMS-DH (aug) 0.82 0.33% (20s) 0.87 2.69% (1.6m) 0.87 0.46% (20s) 0.92 0.78% (1.7m)
GMS-DH Simple 0.76 8.18% (3.3s) 0.83 7.01% (13s) 0.76 12.07% (3.2s) 0.85 8.93% (12s)
H.3 Generalization to Larger Instances
In Tables 11-13, we report more generalization results for our methods in the MOTSP and MGMOTSP cases. Notably, even when introducing significantly more edges by testing on FLEX1010 and FIX1010, our methods remain competitive zero-shot.
Table 11: MOTSP for instances with more nodes.